OpenStudy (anonymous):

solve the equation 11^(6n-11)=1/14641

OpenStudy (anonymous):

if you recognise that 14641=11^4 than it is easy. 1/14641=11^-4 so 6n-11=-4 6n=7 n=7/4

OpenStudy (anonymous):

thats not a choice i have

myininaya (myininaya):

6n=7 solve this for n

OpenStudy (anonymous):

the answer is not 7/4 but 7/6 sorry :-)

myininaya (myininaya):

he just made a typo-0